You drop a ball of mass 49kg  from rest from 6m  above the surface of the Earth. Place the origin of the coordinate system on the surface of the Earth and orient the positive y -direction vertically upward. What is the y -component of the ball's velocity 0.6636s  after being released?
Answer #1

>0 acceleration = -9.81 m/s go Vy = ly tat Ny = 0 -(9.81) (0.6636) > y = -6.51 mis